Find the domain of f (x) = √(x−2)( x − 3)
1
Expert's answer
2019-04-15T11:56:28-0400

"Domain:\;\;(x-2)(x-3)\geq{0} \;\; ->\;\;x\geq{3}\;\;or\;\;x\leq{2}"